# Exciton fractional Chern insulators in moiré heterostructures

### v1

@@ A.2 Exciton–exciton interaction
We extend the method employed to derive the single-particle moiré exciton Hamiltonian [35] to obtain the Hamiltonian describing exciton–exciton interactions, which takes the form
$H_{\text{x-x}}=$ $\frac{1}{2}\sum_{\mathbf{QQ^{\prime}q}}V(\mathbf{q})X^{\dagger}_{\mathbf{Q+q}}X^{\dagger}_{\mathbf{Q^{\prime}-q}}X^{\phantom{\dagger}}_{\mathbf{Q^{\prime}}}X^{\phantom{\dagger}}_{\mathbf{Q}}$
$V(\mathbf{q})=$ $\braket{X_{\mathbf{Q+q}}X_{\mathbf{Q^{\prime}-q}}}{H_{\text{int}}}{X_{\mathbf{Q^{\prime}}}X_{\mathbf{Q}}}.$
For the sake of simplicity, we omit the layer indices and note that the generalization to the multilayer system is straightforward. The matrix element $V(\mathbf{q})$ contains the interaction Hamiltonian in the electron–hole picture,
$H_{\text{int}}=$ $H_{\text{e-e, int}}+H_{\text{h-h, int}}+H_{\text{e-h, int}}$
$H_{\text{e-e, int}}=$ $\frac{1}{2}\sum_{\mathbf{kk^{\prime}q}}V^{\text{e-e}}(\mathbf{q})e^{\dagger}_{\mathbf{k+q}}e^{\dagger}_{\mathbf{k^{\prime}-q}}e^{\phantom{\dagger}}_{\mathbf{k^{\prime}}}e^{\phantom{\dagger}}_{\mathbf{k}}$
$H_{\text{h-h, int}}=$ $\frac{1}{2}\sum_{\mathbf{kk^{\prime}q}}V^{\text{h-h}}(\mathbf{q})h^{\dagger}_{\mathbf{k+q}}h^{\dagger}_{\mathbf{k^{\prime}-q}}h^{\phantom{\dagger}}_{\mathbf{k^{\prime}}}h^{\phantom{\dagger}}_{\mathbf{k}}$
$H_{\text{e-h, int}}=$ $-\sum_{\mathbf{kk^{\prime}q}}V^{\text{e-h}}(\mathbf{q})e^{\dagger}_{\mathbf{k+q}}h^{\dagger}_{\mathbf{k^{\prime}-q}}h^{\phantom{\dagger}}_{\mathbf{k^{\prime}}}e^{\phantom{\dagger}}_{\mathbf{k}},$
The matrix element $V(\mathbf{q})$ thus has contributions from different interaction mechanisms. Here, we focus on the contribution arising from electron–electron interactions, $H_{\text{e-e, int}}$, to illustrate the procedure. First, we expand the exciton states into the electron–hole basis and exploit the fermionic commutation relations together with $e_{\mathbf{k}}\ket{0}=h_{\mathbf{k}}\ket{0}=0$ until all electron and hole operators in the matrix element disappear, obtaining
$\left.V(\mathbf{q})\right|_{\text{e-e}}=\sum_{\mathbf{k}_{i}}\phi^{*}_{\mathbf{k}_{1}}\phi^{*}_{\mathbf{k}_{2}}\phi_{\mathbf{k}_{3}}\phi_{\mathbf{k}_{4}}(A-B)C$
$A=\delta_{-\mathbf{k}_{2}-\alpha_{\text{h}}\mathbf{q},-\mathbf{k}_{3}}\delta_{-\mathbf{k}_{1}+\alpha_{\text{h}}\mathbf{q},-\mathbf{k}_{4}}$
$B=\delta_{-\mathbf{k}_{2}+\alpha_{\text{h}}(\mathbf{Q^{\prime}}-\mathbf{q}),-\mathbf{k}_{4}+\alpha_{\text{h}}\mathbf{Q}}\delta_{-\mathbf{k}_{1}+\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}),-\mathbf{k}_{3}+\alpha_{\text{h}}\mathbf{Q}^{\prime}}$
$C=\braket{0}{e^{\phantom{\dagger}}_{\mathbf{k}_{1}+\alpha_{\text{e}}(\mathbf{Q}+\mathbf{q})}e^{\phantom{\dagger}}_{\mathbf{k}_{2}+\alpha_{\text{e}}(\mathbf{Q}^{\prime}-\mathbf{q})}H_{\text{e-e}}e^{\dagger}_{\mathbf{k}_{3}+\alpha_{\text{e}}\mathbf{Q}^{\prime}}e^{\dagger}_{\mathbf{k}_{4}+\alpha_{\text{e}}\mathbf{Q}}}{0},$
where $i=\{1,2,3,4\}$. Next, we calculate $C$, again by making use of $e_{\mathbf{k}}\ket{0}=h_{\mathbf{k}}\ket{0}=0$ and the commutation relations, and evaluating the Kronecker deltas. The resulting expression reads
$\left.V(\mathbf{q})\right|_{\text{e-e}}=$ $\left.V(\mathbf{q})\right|_{\text{e-e,1}}-\left.V(\mathbf{q})\right|_{\text{e-e,2}}-\left.V(\mathbf{q})\right|_{\text{e-e,3}}+\left.V(\mathbf{q})\right|_{\text{e-e,4}}$
$\left.V(\mathbf{q})\right|_{\text{e-e,1}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}\mathbf{q}}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}\mathbf{q}}V^{\text{e-e}}(\mathbf{q})$
$\left.V(\mathbf{q})\right|_{\text{e-e,2}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}\mathbf{q}}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}\mathbf{q}}$
×$\times V^{\text{e-e}}(\mathbf{k}+\mathbf{q}-\mathbf{k}^{\prime}+\alpha_{\text{e}}(\mathbf{Q}-\mathbf{Q}^{\prime}))$
$\left.V(\mathbf{q})\right|_{\text{e-e,3}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}$
×$\times V^{\text{e-e}}(\mathbf{k}+\mathbf{q}-\mathbf{k}^{\prime}+\alpha_{\text{h}}(\mathbf{Q}-\mathbf{Q}^{\prime}))$
$\left.V(\mathbf{q})\right|_{\text{e-e,4}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}$
×$\times V^{\text{e-e}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime}),$
where each term accounts for direct interaction, electron–electron exchange, hole–hole exchange, and exciton–exciton exchange, respectively [43]. For interlayer excitons, the direct term $V(\mathbf{q})|_{\text{e-e,1}}$ dominates [41] and therefore we disregard the remaining terms. Applying this method to the hole–hole and electron–hole interaction terms yields the exciton–exciton interaction potential in Eq. (3).
Appendix B Additional details of numerical methods
B.1 Many-body Chern number
In order to further assess the topological nature of ground states, we calculate the many-body Chern number, which takes the form
[…]
with $(\Phi_{x},\Phi_{y})$ and |$\ket{\Psi_{i}}$ being the inserted magnetic flux and the many-body ground state $i$, respectively. Here, the integral is evaluated over the $2\pi\times 2\pi$ torus ($T^{2}$). We calculate the Chern number for all many-body ground states and obtain the average many-body Chern number $\mathcal{C}_{\text{avg}}$ for each state [73].
B.2 Finite system geometry
Throughout this work, we consider finite systems (with periodic boundary conditions), the geometry of which is determined by the spanning vectors $\mathbf{T}_{n}\cdot(\mathbf{a_{1}},\mathbf{a_{2}})$, $n=1,2$, with $\mathbf{T}_{1}=(n_{x_{1}},n_{y_{1}})$ and $\mathbf{T}_{2}=(n_{x_{2}},n_{y_{2}})$ and where $\mathbf{a}_{1},\mathbf{a}_{2}$ are the moiré lattice vectors. The system encloses $N_{s}=|n_{x_{1}}n_{y_{2}}-n_{x_{2}}n_{y_{1}}|$ moiré unit cells [74]. To achieve a uniform sampling of the momentum points in the moiré Brillouin zone, we select the spanning vectors for each system size as follows:
  • $N_{\text{s}}=10$: $\mathbf{T}_{1}=(3,1)$, $\mathbf{T}_{2}=(-1,3)$.
  • $N_{\text{s}}=11$: $\mathbf{T}_{1}=(3,-2)$, $\mathbf{T}_{2}=(1,3)$.
  • $N_{\text{s}}=12$: $\mathbf{T}_{1}=(4,0)$, $\mathbf{T}_{2}=(0,3)$.
  • $N_{\text{s}}=13$: $\mathbf{T}_{1}=(3,-2)$, $\mathbf{T}_{2}=(2,3)$.
  • $N_{\text{s}}=14$: $\mathbf{T}_{1}=(4,2)$, $\mathbf{T}_{2}=(-3,2)$.
  • $N_{\text{s}}=16$: $\mathbf{T}_{1}=(4,0)$, $\mathbf{T}_{2}=(0,4)$.
  • $N_{\text{s}}=18$: $\mathbf{T}_{1}=(4,1)$, $\mathbf{T}_{2}=(-2,4)$.
  • $N_{\text{s}}=20$: $\mathbf{T}_{1}=(5,0)$, $\mathbf{T}_{2}=(0,4)$.
Figure 4: Additional data for Abelian states at half filling with long-range interactions. (a) Many-body energy spectrum (10 lowest states) for $N_{\text{s}}=20$ sites. The two degenerate ground states are marked in red, while the blue arrow denotes the energy gap. (b) Scaling of the energy gap including calculations for $N_{\text{s}}=10,12,14,16,18,20$ sites.
Figure 5: Additional data for non-Abelian states at filling one with long range interaction in a system of $N_{\text{x}}=14$ excitons. (a)-(b) Spectral flow for the momentum sectors $K=1$ and $K=8$, respectively, showing that the ground states remain separated from excited states. (c)-(d) PES for the three ground states with $N_{A}=4$ and $N_{A}=5$, respectively. The number of states below the red line in (c) and (d) is 1848 and 4942, respectively, corresponding to the quasi-hole excitation counting for Moore–Read states. The inset in (d) displays a zoomed-in region in the PES showing the sparse distribution of states where the entanglement gap is expected.
Figure 6: Moore–Read states at filling one for systems with $N_{\text{x}}=11$ and $N_{\text{x}}=12$ excitons. Many-body spectrum (10 lowest energies) and PES at filling $\nu=1$ of the exciton Chern band with long-range interactions for (a)-(b) 11 and (c)-(d) 12 particles. The number of states below the red line matches the quasi-hole counting for the Moore–Read state (253 and 328) for 11 and 12 particles with $N_{A}=3$.
Appendix C Supporting numerical evidence
C.1 Abelian states at half filling
The observed signatures for Laughlin states at half filling remain robust across the wide range of system sizes explored. In Fig. 4(a), we show that the two ground states appear in the energy spectra also for the largest system size that is accessible in our numerical calculations, $N_{\text{s}}=20$. These states are present in smaller systems as well. Importantly, the gap with respect to the excited states remains robust ($\sim 10$ meV) for all the studied systems ranging from $N_{\text{s}}=10$ up to $N_{\text{s}}=20$ sites, cf. Fig. 4(b).
C.2 Non-Abelian states at filling one
In the case of Moore–Read states at filling $1$, the gap between ground and excited states for an even number of excitons ($N_{\text{x}}=14$) is small. Nevertheless, we show in Fig. 5(a)-(b) that ground and excited states remain well separated upon threading a magnetic flux. In addition, in Fig. 5(c)-(d) we show that the presence of a PES gap matching the counting of quasi-hole excitations is a robust feature also for a higher number of particles in the $A$ sub-system, $N_{A}=4$, and is still present for $N_{A}=5$, although less pronounced. In Fig. 6 we show that the characteristic features of Moore–Read states are present not only in the largest systems studied ($N_{\text{x}}=13,14$) but also in systems with a lower number of excitons ($N_{\text{x}}=11,12$ is shown here).

### v2

@@ Appendix A Exciton–exciton interaction
We take the method employed for deriving the single-particle moiré exciton Hamiltonian [35] and extend it in order to obtain the Hamiltonian describing exciton–exciton interactions, which takes the form
Here, $\rho_{l}(\mathbf{q})=\sum_{\mathbf{Q}}X^{\dagger}_{l\mathbf{Q+q}}X^{\phantom{\dagger}}_{l\mathbf{Q}}$ is the exciton density operator, :::: denotes normal order, and $V_{ll^{\prime}}(\mathbf{q})$ is the exciton–exciton interaction potential. In the first part of this work, we consider ideal contact interactions of the form $V_{ll^{\prime}}(\mathbf{q})=V_{0}$. Later, we consider the realistic long-range interaction potential, which reads
$V_{ll^{\prime}}(\mathbf{q})=$ $V^{\text{e-e}}_{ll^{\prime}}(\mathbf{q})+V^{\text{h-h}}_{ll^{\prime}}(\mathbf{q})-V^{\text{e-h}}_{ll^{\prime}}(\mathbf{q})-[V^{\text{e-h}}_{l^{\prime}l}(-\mathbf{q})]^{*},$
$V^{\text{e-e}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{h}}\mathbf{q},\alpha_{\text{h}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|},$
$V^{\text{h-h}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{e}}\mathbf{q},\alpha_{\text{e}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|d|l-l^{\prime}|},$
$V^{\text{e-h}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{e}}\mathbf{q},\alpha_{\text{h}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|dl},$ (2)
with $d=0.67$ nm and $\epsilon=3.8$ [83]. The exciton–exciton interaction potential above contains individual contributions accounting for interactions between the charges (electrons and holes) in different excitons, i.e. it includes electron–electron, hole–hole, and electron–hole interaction channels. Here, $F_{ll^{\prime}}(\mathbf{q}_{1},\mathbf{q}_{2})=\mathcal{F}_{ll,\mathbf{q}_{1}}\mathcal{F}_{l^{\prime}l^{\prime},-\mathbf{q}_{2}}$ has been introduced together with the exciton form factor $\mathcal{F}_{ll^{\prime},\mathbf{q}}=\sum_{\mathbf{k}}\phi^{*}_{l\mathbf{k}}\phi_{l^{\prime}\mathbf{k+q}}$, where $\phi_{l\mathbf{k}}$ is the wave function describing the relative electron–hole coordinate within a single exciton. While terms arising from different exchange mechanisms should formally contribute to the interaction, here we have focused on the direct term which should dominate in the context of interlayer excitons [59, 60, 61, 62]. A detailed derivation of the interaction potential can be found in the SM.